y^2=x+1. Find the vertex, x and y intercepts, and sketch the graph. Please help, I have no idea what to do with this!!!
Answer by stanbon(75887)

y^2=x+1. Find the vertex, x and y intercepts, and sketch the graph.
-----------
(y-0)^2 + 4(1/4)(x+1)
------------------
Vertex: (-1,0)
----------------
x-intercept: Let y = 0, then x+1 = 0; x = -1
x-intercept: (-1,0)
-------------------------------
y-intercept: Let x = 0, then y^2 = 1; y = +/-1
y-intercepts: (0,1) and (0,-1)
-------------------------------
Graph:
x = y^2-1
----
Plot the vertex at (-1,0)
Plot the y-intercepts at (0,1) and (0,-1)
The graph is a parabola opening to the right.
=================
Cheers,
Stan H.